OpenStudy (anonymous):

Let V = Z be the set of integers with the usual addition as “vector addition.” Show that it is not possible to define scalar multiplication by Q;R; or C so as to make it into a vector space.

OpenStudy (anonymous):

What IS scalar multiplication as required by vector spaces anyway? :D

OpenStudy (anonymous):

\[(\alpha f)(x) = \alpha f (x) \]

OpenStudy (anonymous):

(f1 + f2)(x) = f1(x) + f2(x)

OpenStudy (anonymous):

Yes, well, some key property of scalar multiplication is off...

OpenStudy (anonymous):

im not sure i understand what you mean

OpenStudy (anonymous):

Well, it's asking you why you can't define scalar multiplication by scalars from Q, R, and C. What is scalar multiplication?

OpenStudy (anonymous):

okay okay i think i get it thank you :)

OpenStudy (anonymous):

Good. :)